OpenStudy (anonymous):

if a container is a glass with a diameter of 2 inches, what is its height?

OpenStudy (anonymous):

No sufficient data given for the problem to be solved

OpenStudy (anonymous):

volume for a cylinder = p(radius)2(height)

OpenStudy (anonymous):

Volume is not given.

OpenStudy (skullpatrol):

$$\Huge\text{ what is its height?}$$

OpenStudy (anonymous):

that is what it is asking for.

OpenStudy (anonymous):

Did I not saw the whole problem? It just said a glass with a diameter of 2? Can you solve for the height if the volume is not given?

OpenStudy (anonymous):

sorry about that. A container holds 16 ounces. what is the volume of this container in inches? i got 28.88 - If the container is a glass with a diameter of 2inches, what is it height?

Directrix (directrix):

Assumption: The glass is a right circular cylinder. V = π * r ² * h 28.88 = π * 1 ² * h @jaycsung Solve for h.